OpenStudy (highschoolmom2010):

Find the value of each variable using the given chord, secant, and tangent lengths. If the answer is not a whole number, round to the nearest tenth.

OpenStudy (highschoolmom2010):



OpenStudy (highschoolmom2010):

(11+20)11=(13+c)13 21*11= 231= that is as far as ive gotten

ganeshie8 (ganeshie8):

right side also distribute and simplify

OpenStudy (highschoolmom2010):

oh ok

OpenStudy (highschoolmom2010):

@ganeshie8 i think that i am doing this wrong (11+20)11=(13+c)13 21*11=169+13c 231= 169+13c -169 -169 62=13c 62/13=13c/13 7.769=c however my book says it is 13.2 i dont see how

ganeshie8 (ganeshie8):

wolfram also says 13.2 let me take a good look at ur work :)

OpenStudy (highschoolmom2010):

ok

ganeshie8 (ganeshie8):

mistake in second line

ganeshie8 (ganeshie8):

(11+20)11=(13+c)13 21*11=169+13c ******************** 231= 169+13c -169 -169 62=13c 62/13=13c/13 7.769=c

ganeshie8 (ganeshie8):

11+20 = 31 not 21 ok ?

OpenStudy (highschoolmom2010):

oh hehe i cant add now -_-

OpenStudy (highschoolmom2010):

thanks

OpenStudy (highschoolmom2010):

i got 13.2 now

ganeshie8 (ganeshie8):

lol

ganeshie8 (ganeshie8):

good :)